The mean income per person in the United States is $35,500, and the distribution of incomes follows a normal distribution. A random sample of 8 residents of Wilmington, Delaware, had a mean of $39,500 with a standard deviation of $8,200. At the 0.010 level of significance, is that enough evidence to conclude that residents of Wilmington, Delaware, have more income than the national average?

  1. State the null hypothesis and the alternate hypothesis.
  1. State the decision rule for 0.010 significance level. (Round your answer to 3 decimal places.)
  1. Compute the value of the test statistic. (Round your answer to 2 decimal places.)
  1. Is there enough evidence to substantiate that residents of Wilmington, Delaware, have more income than the national average at the 0.010 significance level?
Answer #1

Solution:

Given: The mean income per person in the United States is $35,500, and the distribution of incomes follows a normal distribution.

\mu = 35500

Sample size = n = 8

Sample mean = \bar{x} =39500

Sample standard deviation = s = 8200

level of significance = 0.01

We have to test if there is evidence to conclude that residents of Wilmington, Delaware, have more income than the national average.

Part a) State the null hypothesis and the alternate hypothesis.

H_{0}: \mu = 35500

H_{1}: \mu > 35500

Part b) State the decision rule for 0.010 significance level.

Since sample size n is small and population standard deviation is unknown and population is Normally distributed, we use t test.

thus find t critical value

df = n - 1 = 8 -1 = 7

One tail area = 0.01

t critical value = 2.998

Thus reject null hypothesis H0 , if t test statistic value \geq 2.998

Part c) Compute the value of the test statistic.

t=\frac{\bar{x}-\mu }{s /\sqrt{n}}

t=\frac{39500-35500 }{8200 /\sqrt{8}}

t=\frac{4000 }{8200 /2.828427 }

t=\frac{4000 }{2899.1378 }

t= 1.380

Part d) there enough evidence to substantiate that residents of Wilmington, Delaware, have more income than the national average at the 0.010 significance level?

Since t= 1.380 < t critical value = 2.998, we fail to reject H0,

thus there is not enough evidence to substantiate that residents of Wilmington, Delaware, have more income than the national average at the 0.010 significance level